OpenStudy (anonymous):

Suppose that a sequence (an) satisfies the condition |an+1 − an+2| < 1/2 |an − an+1|. Prove that (an) converges.

OpenStudy (anonymous):

i think this is hard, and may need induction you want to show the sequence is a cauchy sequence

OpenStudy (anonymous):

since \(|a_3-a_2|<\frac{1}{2}|a_2-a_1|\) we know \[|a_4-a_3|<\frac{1}{2}|a_3-a_2|<\frac{1}{2^2}|a_2-a_1|\]

OpenStudy (anonymous):

by induction you get \[|a_{n+2}-a_{n+1}|<\frac{1}{2^n}|a_2-a_1|\] which is getting closer to proving it is a cauchy sequence, but not quite there yet

OpenStudy (anonymous):

perhaps you had an example in class where you had to show if \[|a_{n+1}-a_n|<2^{-n}\] then the sequence was a cauchy sequence you can mimic that here

OpenStudy (anonymous):

What if I didn't

OpenStudy (anonymous):

hmm then you have some work to do depends on what you can use it is a fact that if \(|a_{n+1}-a_n|<b_n\) where \(\sum b_n\) is finite, then \({a_n}\) is cauchy but if you did not do this , then what you have to show is that \[|a_n-a_m|\to 0\] for arbitrary \(m,n\) in this case you will need the triangle inequality repeatedly

OpenStudy (anonymous):

best bet is to google the proof that if \(|a_{n+1}-a_n|<\frac{1}{n^2}\) then \(\{a_n\}\) is cauchy

OpenStudy (anonymous):

look at #4 here http://math.stackexchange.com/questions/182830/proving-that-a-sequence-is-cauchy more than i can write out, but it is all there

OpenStudy (anonymous):

ok that was a typo, i meant google the proof that \(|a_{n+1}-a_n|<\frac{1}{2^n}\) etc i reversed the \(n\) and \(2\) by mistake

OpenStudy (anonymous):

Thank you kind sir. May God be with you in your travels